a 100-m rope is cut into 2 pieces.one is 16m greater than the other.find the lenght of eah piece
Answer by emargo19(101)

Let one piece of the rope be X
Let the other piece of the rope be x+16
Equation is :
Piece 1 + Piece 2 = 100
x + (x+16) = 100
2x = 100-16
2x= 84
x= 42
Piece 1(x)= 42 m.
Piece 2 (x+16)= 42+16=58 m.

check :
42+ 58 = 100